Why Iron has electron shel 2,8,14,2 and why it has valency 2+ instead of 2-

Grade:7

1 Answers

Ansh Agrawal
7 years ago
Electron configurations of elements can be a bit confusing, and the transition metals are a case in point.  Iron has the configuration 1s^2 2s^2 2p^6 3s^2 3p^6 4s^2 3d^6.  Note that the 4s orbitals are filled before you add six electrons to the 3d orbitals.  Yet, the 3d orbitals do actually belong with the n=3 orbitals and not the n=4 orbitals.  So, you will have 2 electrons in the n=1 (1s) shell, 8 in the n=2 (2s and 2p) shell, (8+6 =) 14 in the n=3 (3s, 3p, and 3d) shell, and finally 2 electrons in the n=4 (4s) shell.  

If, perhaps, this wasn't enough fun for you, take a look at the electronic configuration of Neodymium (Nd) and other Lanthanides.  These have incompletely filled n=4, n=5, and n=6 shells.